### v1

# Untying the Reversal Curse via Bidirectional Language Model Editing

Abstract
Recent studies have demonstrated that large language models (LLMs) store massive factual knowledge within their parameters. But existing LLMs are prone to hallucinate unintended text due to false or outdated knowledge. Since retraining LLMs is resource intensive, there has been a growing interest in the concept of model editing. Despite the emergence of benchmarks and approaches, these unidirectional editing and evaluation have failed to explore the reversal curse. Intuitively, if “The capital of France is” is edited to be a counterfact “London” within a model, then it should be able to naturally reason and recall the reverse fact, i.e., “London is the capital of” followed by “France” instead of “England”. In this paper, we study bidirectional language model editing, aiming to provide rigorous model editing evaluation to assess if edited LLMs can recall the editing knowledge bidirectionally. A new evaluation metric of reversibility is introduced and a benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed to evaluate the reversibility of edited models in recalling knowledge in the reverse direction of editing. We surprisingly observe that while current editing methods and LLMs can effectively recall editing facts in the direction of editing, they suffer serious deficiencies when evaluated in the reverse direction. To mitigate the reversal curse, a method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed. A set of editing objectives that incorporate bidirectional relationships between subject and object into the updated model weights are designed. Experiments show that BIRD improves the performance of four representative LLMs of different sizes via question answering and judgement.
Machine Learning, ICML
Figure 1: Comparison of existing unidirectional evaluation paradigms and the proposed paradigm that assesses model editing bidirectionally. Assessing whether edited models can recall the editing facts (a) via single-hop questions in the direction of editing; (b) via multi-hop questions in the direction of editing; (c) in the reverse direction of editing. $f_{\theta}$ / $f_{\theta_{e}}$ denotes the models before / after editing.
1 Introduction
Large language models (LLMs) have shown impressive capabilities in understanding and generating text, improving performance on a variety of natural language processing tasks such as code generation (Chen et al., 2021; Mitchell et al., 2022a), open-ended dialogue (Wu et al., 2023; Ouyang et al., 2022) and machine translation (Liu et al., 2022; Zhang et al., 2023a). However, existing LLMs inevitably exhibit hallucinations due to false or outdated knowledge stored in their parameters (Zhang et al., 2023b; Peng et al., 2023; Ji et al., 2023), which degrades the system performance and fails to meet user expectations in many real-world scenarios. Considering the huge computing resource consumption of retraining LLMs and echoing green deep learning (Schwartz et al., 2020; Xu et al., 2021), there has been considerable attention on the research regarding model editing (Sinitsin et al., 2020; Zhu et al., 2020; Cao et al., 2021; Meng et al., 2022, 2023; Dai et al., 2022; Mitchell et al., 2022b; Yao et al., 2023), which aims at modifying knowledge by updating model weights.
Existing methods developed to modify the parameters of models can be generally categorized into two paradigms (Wang et al., 2023; Yao et al., 2023). (1) Meta-learning: these methods employ an editor hypernetwork to learn the adjustments ()($\triangle$)( △ ) for parameter updates in the base model (Cao et al., 2021; Mitchell et al., 2022a). (2) Locate-Then-Edit: these approaches locate knowledge neurons in LLMs that are positively correlated to a knowledge expression and then apply modifications to them accordingly (Dai et al., 2022; Meng et al., 2022, 2023). Currently, the effectiveness of an editing method is usually evaluated along three dimensions. (1) Efficacy is to validate whether the post-editing model can recall the exact editing fact. (2) Generalization is to verify whether the edited model can recall the editing fact under different paraphrase prompts. (3) Locality [^1] Also referred to as specificity in Meng et al. (2022, 2023). is to verify whether the output of the edited model for unrelated inputs remains unchanged after editing. Recently, researchers have made strides in model editing for LLMs in various editing tasks and settings. Both the ZsRE (Levy et al., 2017) and CounterFact (Meng et al., 2022) datasets utilize the aforementioned evaluation metrics to assess whether edited models can recall the newly injected facts as depicted in Figure 1(a), and remain unaffected for unrelated knowledge. To further assess the model editing capabilities, Yao et al. (2023) have taken a step forward to introduce an evaluation metric regarding portability, which is designed to gauge the effectiveness of model editing in transferring knowledge to related content. Besides, MQuAKE dataset (Zhong et al., 2023) is constructed to extend the assessment of generalization by measuring whether edited models correctly answer the multi-hop questions entailed by the editing facts, as shown in Figure 1(b). Despite the emergence of benchmarks and techniques, these editing and evaluation approaches operate under the unidirectional paradigm following only the direction being edited. The reversal curse (Berglund et al., 2023) has not yet been explored in model editing, and it remains unclear whether edited models can reason and recall the editing facts in the reverse direction of editing, as shown in Figure 1(c).
In this paper, we study bidirectional language model editing, with the goal of conducting a comprehensive evaluation of model editing to determine if edited LLMs can accurately recall the editing knowledge bidirectionally. Among various kinds of learned beliefs such as logical, spatial, or numerical knowledge, this paper investigates factual knowledge within LLMs [^2] Model editing is also termed as knowledge editing.. In addition to the metrics of efficacy, generalization and locality that are commonly used in previous works, a new evaluation metric of reversibility is introduced to evaluate the effectiveness of edited models in recalling knowledge in the reverse direction of editing. To evaluate the reversibility of edited models, a benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed, where each piece of knowledge is represented in the form of a triple (subject, relation, object). When inverting the relation between subject and object, the taxonomy designed in Bordes et al. (2013) which categorizes relations between entities into four classes: one-to-one, one-to-many, many-to-one, many-to-many, is considered. Furthermore, two evaluation forms of question answering (QA) and judgment are adopted to characterize these symmetrical and asymmetrical relations.
To establish the performance comparison of existing model editing methods on bidirectional model editing, this paper presents meticulous experiments involving as many methods as possible such as FT (Zhu et al., 2020), MEND (Mitchell et al., 2022a), KN (Dai et al., 2022), ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023). Four representative LLMs of different sizes are selected as the base models: GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang & Komatsuzaki, 2021), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). We surprisingly observe that while current model editing methods and LLMs perform significantly well at recalling the editing facts in the editing direction, they suffer serious deficiencies when evaluated in the reverse direction. For instance, though a LLaMA-2 model edited by the state-of-the-art method ROME (Meng et al., 2022) could recall 99.70% editing facts in the editing direction, it can only recall 0.26% editing facts in the reverse direction.
Towards bidirectional language model editing, a simple yet effective method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed to enhance the effectiveness of edited models in recalling the editing facts in the reverse direction. A set of editing objectives are designed to incorporate the bidirectional and invertible relationships between subject and object in a given fact into the updated model weights. Inspired by Mikolov et al. (2013), the representation of the newly editing object is driven close to the representation of the subject plus forward relation. Reversely, the representation of the subject is also driven close to the representation of the editing object plus the backward relation. Here, the relationships along both directions are inversible to each other. By this means, it can help characterize the bidirectional and inversible relationships between subject and object in factual knowledge, enabling models to establish factual associations bidirectionally during the editing process. Moreover, the proposed editing method eliminates the training step and does not require extra training data. Experiments with four representative LLMs of different sizes demonstrate that the proposed editing method not only improves the generalization ability when recalling the editing facts in the direction of editing, but also achieves highly promising performance in the reverse direction.
In summary, our contributions are three-fold: (1) This study makes the first attempt to explore bidirectional language model editing and evaluation. A new evaluation metric of reversibility is introduced and a new benchmark BAKE is constructed to assess the reversibility of editing methods. (2) This paper presents meticulous experiments on existing popular editing methods and LLMs. Through analysis, we surprisingly found that existing methods and LLMs suffer serious deficiencies when evaluated in the reverse direction of editing. (3) A simple yet effective method BIRD is proposed to incorporate the bidirectional and inversible relationships between subject and object in a given fact into the updated model weights, which mitigate the reversal curse in model editing. To facilitate others to reproduce our results, we will publish all data and source code later at https://github.com/mjy1111/BAKE. We hope that both our benchmark and method can shed light on bidirectional language model editing.
2 Related Work
Model Editing
Many studies have investigated model editing, including memory-based, meta-learning, and locate-then-edit (Yao et al., 2023). Memory-based methods do not modify model weights but store the editing facts with an external memory (Mitchell et al., 2022b; Zhong et al., 2023). For example, Mitchell et al. (2022b) stored edits in a base model and learned to reason over them to adjust its predictions as needed. The latter two classes of methods are developed to directly modify the internal parameters of models, which is the focus of this paper. On the one hand, meta-learning methods train a hypernetwork to get gradient changes to update model parameters (Cao et al., 2021; Mitchell et al., 2022a). Cao et al. (2021) utilized a hypernetwork to predict parameter shift at test time. Mitchell et al. (2022a) learned to transform the fine-tuning gradient into a low-rank decomposition of the gradient. On the other hand, locate-then-edit methods first locate knowledge neurons in LLMs that exhibit a positive correlation with a knowledge expression, and then modify them accordingly (Dai et al., 2022; Meng et al., 2022, 2023). In particular, Dai et al. (2022) computed the contribution of each neurons to a certain knowledge, then updated or erased knowledge by modifying these neurons with the embedding vectors of facts. Meng et al. (2022) located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair in the MLP module.
The Reversal Curse of LLMs
Recent works have shown a surprising failure of generalization in auto-regressive LLMs. Berglund et al. (2023) have verified that a model fine-tuned on a sentence of the form “A is B”, it will not automatically generalize to the reverse direction “B is A”. Moreover, the likelihood of the correct answer will not be higher than for a random answer. This is termed as the reversal curse. Contemporary to this work, Grosse et al. (2023) used influence functions to determine how much adding a given training example influences an LLM’s outputs. For instance, given the input A, what most influences the likelihood of B? In their experiments, training examples that match the order (“A precedes B”) are far more influential than examples with reverse order (“B precedes A”). Results show that if a pretrained model was not trained on facts in both directions, it would not generalize to both directions.
Interpretability of Transformer
Many studies have attempted to explain and understand Transformer (Vaswani et al., 2017). Previous works focus on the function of self-attention layers and the differences between layers (Tenney et al., 2019; Vig & Belinkov, 2019; Voita et al., 2019; Clark et al., 2019). For instance, Voita et al. (2019) found that the most confident attention heads played consistent linguistically-interpretable roles. More recently, a surge of works have investigated the role of feed-forward network (FFN) layers in Transformer, showing that FFN layers store factual and linguistic knowledge, which can be located in specific neurons and edited (Geva et al., 2021; Da et al., 2021; Meng et al., 2022; Dai et al., 2022). Besides, Geva et al. (2022) investigated the mechanism in which FFN layers update the inner representations of Transformer-based LMs. They propose that the FFN output can be viewed as a collection of updates that promote concrete concepts in the vocabulary space.
Compared with previous studies (Dai et al., 2022; Meng et al., 2022, 2023; Mitchell et al., 2022a; Yao et al., 2023; Zhong et al., 2023) that are the most relevant to our work, a main difference should be highlighted. These approaches target only on unidirectional editing and evaluation of model editing, while this study explores bidirectional editing and evaluation. To the best of our knowledge, this paper makes the first attempt to introduce a new evaluation metric of reversibility, build a benchmark for evaluating the reversibility of editing methods, and propose a simple yet effective method to mitigate the reversal curse in model editing.
3 Preliminary
This section introduces the definition of model editing and the preliminary knowledge of a popular editing method ROME (Meng et al., 2022).
3.1 Problem Definition
The goal of model editing is to insert new facts into model parameters without retraining. In this paper, we study facts of the form ($s,r,o)$, consisting of a subject $s$, a relation $r$, and an object $o$ (e.g., $s$ = Eiffel Tower, $r$ = locate in, $o$ = Paris). Editing a fact is to insert a new knowledge triple $(s,r,o^{*})$ in place of the current triple $(s,r,o)$, where these two triples share the same subject and relation. An editing operation is represented as $e=(s,r,o,o^{*})$ for brevity.
Given a set of editing facts =$\mathcal{E}=\left\{e_{1},e_{2},\ldots\right\}$ and a model $f$ , model editing involves learning a function $K$ that yields an edited language model $f^{*}:K(f,\mathcal{E})=f^{*}$. To evaluate the effectiveness of various model editing methods, previous works focus on evaluation along three dimensions: efficacy, generalization and locality (Cao et al., 2021; Mitchell et al., 2022a; Meng et al., 2022).
Efficacy
Given an editing fact ($s$ = Eiffel Tower, $r$ = located in, $o$ = Paris, $o^{*}$=London), it could be regarded edited effectively if the edited model $f^{*}$ assigns a higher probability to the statement “Eiffel Tower is located in London” than the original prediction (Paris).
Generalization
Edited models should be able to recall the updating knowledge when prompted within the editing scope. The paraphrase prompts of the question like “Where is the Eiffel Tower?” or “What is the location of the Eiffel Tower?” can be used. The edited model $f^{*}$ generalizes successfully if it could recall the editing memory “London”.
Locality
The edited model $f^{*}$ should remain unchanged on the irrelevant examples out of the editing scope. For example, the answer to “Where is the Louvre?” should still be “Paris”.
3.2 Rank-One Model Editing (ROME)
ROME (Meng et al., 2022) applies causal mediation analysis (Pearl, 2001; Vig et al., 2020) to locate the MLP modules that store facts. An MLP module consists of two layers, where the first and second layers are denoted as $W_{fc}^{l}$ and $W_{proj}^{l}$ respectively. $W_{proj}^{l}$ is considered as a linear associative memory and an editing area. A new fact is represented as a key-value pair ($k_{*}$, $v_{*}$), which can be inserted into the MLP module of the model by solving a constrained least-squares problem (Bau et al., 2020). All that remains is to choose the appropriate $k_{*}$ and $v_{*}$.
Denote $L$ as an edited MLP layer, $G$ as the model, and $e=(s,r,o,o^{*})$ as an editing operation. The key $k_{*}$ is obtained by calculating the average representation of the last token of the subject $s$ outputted by $W_{fc}^{L}$ as follows:
$k_{*}$ $=\frac{1}{N}\sum_{j=1}^{N}k\left(x_{j}+s\right)\text{,}$ (1)
where $\text{where }k(x)$ $=\sigma\left(W_{fc}^{\left(L\right)}\gamma\left(a_{[x],i}^{\left(L\right)}+h_{[x],i}^{\left(L-1\right)}\right)\right)\text{. }$
Here, $x_{j}$ is the random sampled sequence and $i$ is the location of the last token of the subject in sentence $x$. $h$ is the hidden states, $a$ is the attention, $\gamma$ is the layernorm, and $\sigma$ is the activation function.
The value $v_{*}$ encodes the new knowledge ($r$, $o^{*}$) as a property of $s$. To calculate it, ROME sets $v_{*}=\operatorname{argmin}_{z}\mathcal{L}(z)$, where the objective is denoted as:
$\mathcal{L}(z)=\frac{1}{N}\sum_{j=1}^{N}-\log\mathbb{P}_{G\left(m_{i}^{\left(L\right)}:=z\right)}\left[o^{*}\mid x_{j}+p\right]+$ (2)
$D_{\mathrm{KL}}\left(\mathbb{P}_{G\left(m_{i^{\prime}}^{\left(L\right)}:=z\right)}\left[x\mid p^{\prime}\right]\|\mathbb{P}_{G}\left[x\mid p^{\prime}\right]\right)\text{. }$
The first term in Eq. (2) seeks a vector z that, when substituted as the output of the MLP at the token $i$ (notated $G\left(m_{i}^{\left(L\right)}:=z\right)$ and here $i$ is the end of the subject in prompt $x_{j}+p$), will cause the network to predict the target object $o^{*}$ in response to the factual prompt $p$ (e.g., $p$=Eiffel Tower is located in, $o^{*}$=London). The second term minimizes the KL divergence (Ma et al., 2022) of predictions for the prompt $p^{\prime}$ (of the form ” { subject } is a”) to the unchanged model, which helps preserve the model’s understanding of the subject’s essence ($i^{\prime}$ is the location of the last token of the subject in $p^{\prime}$).
Finally, the MLP weight $W_{proj}^{L}$ is updated with a rank-one update to insert the new fact:
$\hat{W}=W+\Lambda\left(C^{-1}k_{*}\right)^{T}\text{,}$ (3)
where $C$ is a constant by estimating the covariance of many existing keys. $\Lambda=\left(v_{*}-Wk_{*}\right)/\left(C^{-1}k_{*}\right)^{T}k_{*}$, representing the residual error of the new key–value pair on the original memory matrix.
Readers can refer to Meng et al. (2022) for more details of ROME.
4 BAKE: Bidirectional Assessment for Knowledge Editing
The BAKE benchmark comprises two datasets of BAKE-Q&J and BAKE-J. Both datasets are designed for evaluating counterfactual edits in LLMs following Meng et al. (2022) and Zhong et al. (2023). When inverting the relation between subject and object, two evaluation forms of question answering (Q) and judgment (J) are adopted to characterize these symmetrical and asymmetrical relations.
4.1 Relation Category
There are various relations between entities in factual knowledge, which could be categorized into four classes: one-to-one, one-to-many, many-to-one, and many-to-many (Bordes et al., 2013), as shown in Figure 2.
For one-to-one and one-to-many relations, both the question answering and judgment forms are employed, since there is only one definite output after inverting the relation. Take ”developed” relation from one-to-many category in Figure 2(b) as an example, one can construct the factual statement “Apple developed iPhone”. When it is validated reversely, it could be presented to models in QA form like “iPhone is developed by ” or judgement form like “Whether iPhone is developed by Apple?” Similarly, the one-to-one category shares the same property as the one-to-many category.
On the other hand, only the judgment form is employed for many-to-one and many-to-many relations, since there are alternative outputs to a given reversal relation. Suppose a “written by” relation from many-to-one category shown in Figure 2(c) is chosen to form a statement “Hamlet is written by William Shakespeare”. If presented to models for reverse validation in QA form like “William Shakespeare has written ”, there may be many correct answers such as King Lear and Macbeth. It is notable that edited models are expected to output Hamlet intended for reversal evaluation instead of King Lear and Macbeth. Therefore, evaluating the reversal relation for many-to-one and many-to-many categories via judgment is better. Similarly, the many-to-many category shares the same property as the many-to-one category.
Figure 2: Examples of four relation categories. A given relation is considered (a) one-to-one if a subject can be associated with at most one object; (b) one-to-many if a subject can be associated with multiple objects; (c) many-to-one if multiple subjects can be linked to the same object; or (d) many-to-many if multiple subjects can be linked to multiple objects.
4.2 Data Construction of BAKE-Q&J
This dataset is constructed based on Wikidata (Vrandecic & Krötzsch, 2014), a knowledge base containing millions of fact triples. First, we manually chose a total of 25 one-to-one and one-to-many relations, and all fact triples under each relation, denoted as $F(r)$= {$(s,r,o)$mid $s$ and $o$ are linked by $r$ in Wikidata}. For each relation $r$, some handwritten templates $T(r)$ and those $T(r^{-1})$ corresponding to its inverse relation $r^{-1}$ were designed [^3] The inverse relation $r^{-1}$ and its templates only adopted in evaluation, not for editing.. Readers can refer to Appendix A for the examples of these templates. Entities can be replaced to form full prompts: $\mathcal{P}(s,r)$ = {$t$.format($s$) $\mid$ $t\in T(r)$}, where .format() is string substitution. For example, a template for (r = written by) might be “{} is written by”, where “Hamlet” substitutes “{}”.
To construct an editing example, two triples sharing the relation $r$ are sampled from Wikidata, denoted as $(s,r,o)$ and $(s^{*},r,o^{*})$ respectively. Following Meng et al. (2022) and Zhong et al. (2023), counterfactual edits were built and evaluated in this paper.
An edit is represented as =$\mathcal{E}=\left\{s,r,o,o^{*},p\right\}$ to test efficacy, where the editing prompt $p\sim\mathcal{P}(s,r)$. Besides, to test generalization, a set of two semantically-equivalent paraphrase prompts $\mathcal{P}^{G}$ are sampled from $\mathcal{P}(s,r)\backslash\left\{p\right\}$. Moreover, to test locality, a set of triples that share the same object $o$ are collected: $\mathcal{S}=\left\{\left(s^{\prime},r^{\prime},o\right)\right\}$. For example, select ($s$= Eiffel Tower, $r$= located in, $o$= Paris), $\mathcal{S}$ might contain triple like ($s$= France, $r$= has capital, $o$= Paris). Then a set of prompts $\left\{\mathcal{P}\left(s^{\prime},r^{\prime}\right)\mid(s^{\prime},r^{\prime},o)\in\mathcal{S}\right\}$ are constructed to sample the neighborhood prompts $\mathcal{P}^{L}$.
As for the reversibility proposed in this paper, after modifying the knowledge from $(s,r,o)$ to $(s,r,o^{*})$, the edited model should be capable of deducing $(o^{*},r^{-1},s)$ if it is reversible. Following Yao et al. (2023) and  Zhong et al. (2023), we should ensure the triple $(o^{*},r^{-1},s)$ is not included (editing knowledge should not pre-exist) and $(o^{*},r^{-1},s^{*})$ is included (original knowledge should pre-exist) in the base LLM. If not, this editing example will be filtered out accordingly. Specifically, given $o^{*}$ and $r^{-1}$, we tested if a model can output $s^{*}$ rather than $s$, to deduce if the triple $(o^{*},r^{-1},s^{*})$ is included in the model. Furthermore, there are QA and judgment forms in this dataset and the reverse-qa prompt is represented as $\mathcal{P}^{Rq}=\left\{o^{*},r^{-1},s^{*},s,p^{-1}\right\}$, where $p^{-1}\sim\mathcal{P}(o^{*},r^{-1})$ and $s$ is the desired answer after editing (e.g., if $s$=Apple, $s^{*}$=Microsoft, $r$=developed, $o^{*}$=Windows, then $r^{-1}$=developed by, and $p^{-1}$ might be “Windows is developed by”). Besides, the reverse-judge prompt $\mathcal{P}^{Rj}$ is a judgment question based on $p^{-1}$, for instance, “Whether Windows is developed by Apple?”, which should be answered with “yes” after editing or “no” before editing.
Table 1: An example in the BAKE-Q&J dataset. $a$ and $a^{*}$ in different columns represent the original answer and desired answer after editing respectively.
$\mathcal{E}$ $(s,r,o)$: (Paris, capital of, France)
$(s^{*},r,o^{*})$: (London, capital of, England)
(Paris, capital of, France $\rightarrow$ England)
$p$: Paris is the capital of ($a$: France, $a^{*}$: England)
$\mathcal{P}^{G}$ Paris, which is the capital of
Paris is the political center of which country?
($a$: France, $a^{*}$: England)
$\mathcal{P}^{L}$ Palace of Versailles and Louvre Museum are located in
Marseille is the city of
Victor Hugo is a famous writer born in
($a$: France, $a^{*}$: France)
$\mathcal{P}^{Rq}$ The capital of England is ($a$: London, $a^{*}$: Paris)
$\mathcal{P}^{Rj}$ Whether the capital of England is Paris? ($a$: no, $a^{*}$: yes)
4.3 Data Construction of BAKE-J
The construction of this dataset is similar to BAKE-Q&J, and the difference lies in that 20 many-to-one and many-to-many relations in total were manually selected. Moreover, only the judgment form is adopted for validating the reversibility, which has been explained in Section 4.1.
Table 2: The statistics of both BAKE-Q&J and BAKE-J.
Type $Num_{BAKE-Q\&J}$ $Num_{BAKE-J}$
Editing examples 4,694 3,184
Relations 25 20
Paraphrase prompts 8,811 5,917
Neighborhood prompts 16,378 23,770
Reverse-qa prompts 4,694 -
Reverse-judge prompts 4,694 3,184
4.4 Dataset Summary
Dataset Format
As shown in Table 1, each example in the BAKE benchmark is represented as a tuple ($\mathcal{E}$, $\mathcal{P}^{G}$, $\mathcal{P}^{L}$,$\mathcal{P}^{Rq}$,$\mathcal{P}^{Rj}$), where $\mathcal{E}$ is the editing knowledge which will be injected into the language model. $\mathcal{P}^{G}$ and $\mathcal{P}^{L}$ are the prompts utilized to validate generalization and locality respectively. $\mathcal{P}^{Rq}$ and $\mathcal{P}^{Rj}$ correspondingly represent the reverse-qa and reverse-judge prompts adopted to evaluate reversibility. $a$ and $a^{*}$ denote the original answer and desired answer after editing respectively.
Dataset Statistics
Table 2 summarizes the statistics of the BAKE-Q&J and BAKE-J datasets. To verify generalization and locality, there is at least one prompt for each editing example. Meanwhile, there is one reverse-qa prompt (only for BAKE-Q&J) and one reverse-judge prompt to verify reversibility. Generally speaking, both datasets consist of counterfactual edits and are utilized to investigate the ability of edited models to recall editing facts in both the editing direction and the reverse direction.
Table 3: Evaluation results (%) of the BAKE-Q&J dataset. The settings for these methods were the same with the published literature. The row named after the model name represents the results evaluated on the base model. Numbers marked with $\dagger$ denoted that the improvement over the best performing baseline was statistically significant (t-test with $p$-value <$0.05$).
Editor Score Efficacy Generalization Locality Reversibility
$\mathrm{S}\uparrow$ $\mathrm{ES}\uparrow$ $\mathrm{EM}\uparrow$ $\mathrm{PS}\uparrow$ $\mathrm{PM}\uparrow$ $\mathrm{NS}\uparrow$ $\mathrm{NM}\uparrow$ $\mathrm{RQS}\uparrow$ $\mathrm{RQM}\uparrow$ $\mathrm{RJS}\uparrow$ $\mathrm{RJM}\uparrow$
GPT-2 XL (1.5B) 0.00 29.12 -4.33 29.57 -4.13 70.15 3.93 0.00 -12.10 0.00 -0.04
FT 20.09 77.38 13.21 50.75 0.40 63.23 2.58 2.82 -11.14 10.46 -0.03
KN 27.41 30.77 -2.52 29.7 -2.64 70.68 2.46 14.42 -6.79 16.05 -0.03
MEND 3.89 93.29 29.85 77.89 9.64 27.82 -7.16 1.82 -12.27 0.25 -0.03
MEMIT 16.54 73.43 20.87 62.77 7.39 64.31 1.89 0.29 -12.19 9.88 -0.02
ROME 41.55 98.74 78.21 91.05 35.53 52.34 -3.54 3.44 -10.13 32.25 -0.01
BIRD $\textbf{42.68}^{\dagger}$ 98.7 77.75 91.32 25.32 52.19 -3.42 2.91 -10.1 $\textbf{34.50}^{\dagger}$ 0.00
GPT-J (6B) 0.00 22.92 -8.34 23.00 -8.36 77.15 8.01 0.00 -19.58 0.00 -0.73
FT 12.70 73.11 11.67 40.59 -3.14 72.82 6.69 0.68 -19.75 6.90 -0.65
KN 41.77 38.99 -1.30 38.57 -1.35 61.09 1.34 41.08 -3.43 30.81 -0.03
MEND 2.36 94.29 43.95 72.50 13.19 35.91 -6.88 0.87 -20.45 0.35 -0.75
MEMIT 21.73 96.21 84.69 86.99 43.19 64.59 -0.24 0.16 -19.82 13.47 -0.62
ROME 42.53 98.01 85.63 94.60 59.75 53.27 -8.70 2.80 -16.54 33.89 -0.25
BIRD 43.25 99.33 88.49 96.00 62.20 52.14 -10.01 2.90 -16.45 $\textbf{34.96}^{\dagger}$0.23
LLaMA-1 (7B) 0.00 19.41 -21.31 19.99 -12.33 80.96 12.22 0.00 -25.73 0.00 -2.00
FT 13.50 99.28 25.01 91.86 19.21 30.45 -7.33 0.20 -24.00 8.05 -1.62
KN 29.23 28.21 -4.86 29.06 -6.42 75.35 8.73 10.58 -20.06 26.67 -1.09
MEND 5.12 75.00 16.13 60.62 2.66 47.23 1.64 0.46 -27.93 2.28 -1.94
MEMIT 40.95 99.28 91.81 95.34 55.03 58.38 -4.93 0.15 -26.01 33.19 -0.73
ROME 41.79 99.19 92.92 94.13 42.52 62.11 0.40 0.20 -25.73 33.75 -0.69
BIRD $\textbf{43.61}^{\dagger}$ 99.33 93.42 95.67 44.9 61.95 1.87 0.10 -25.75 $\textbf{36.22}^{\dagger}$0.59
LLaMA-2 (7B) 0.00 16.94 -16.57 17.09 -16.60 83.35 16.55 0.00 -34.00 0.00 -0.32
FT 13.78 97.72 23.43 91.37 19.47 32.66 -6.92 0.21 -32.13 8.18 -0.35
KN 38.50 46.57 -8.27 45.14 -9.68 50.61 10.81 12.70 -24.72 36.67 0.16
MEND 16.39 91.10 49.75 70.27 16.83 40.89 -5.32 0.48 -33.48 9.81 -0.29
MEMIT 38.92 99.85 91.76 96.92 54.98 61.33 -1.73 0.14 -34.20 30.10 -0.23
ROME 47.39 99.70 96.44 94.74 50.93 66.30 3.35 0.26 -33.86 40.78 -0.10
BIRD $\textbf{48.40}^{\dagger}$ 99.90 96.20 94.76 50.44 64.25 3.11 0.28 -33.85 $\textbf{42.72}^{\dagger}$0.07
5 Experiments
In this section, experiments were conducted to demonstrate the reversal curse of existing model editing methods based on four representative LLMs of different sizes.
5.1 Base LLMs and Editing Methods
Given limited computing resources, experiments were conducted on four LLMs of different sizes including GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang & Komatsuzaki, 2021), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). Readers can refer to the corresponding papers of each LLM for details. Five popular model editing methods were selected as the baselines including FT (Zhu et al., 2020), KN (Dai et al., 2022), MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023). Readers can refer to Appendix B.1 for the details of these editing methods.
5.2 Evaluation Metrics
In addition to the proposed Reversibility, three basic metrics of Efficacy, Generalization and Locality (Meng et al., 2022, 2023), as well as two follow-up metrics of Portability (Yao et al., 2023) and Multi-hop (Zhong et al., 2023) were adopted to comprehensive measure each editing method. Given two facts of $(s,r,o)$ and $(s^{*},r,o^{*})$, one editing operation was to update $(s,r,o)$ to $(s,r,o^{*})$. Each metric was calculated as follows.
  • $\bullet$
    Efficacy validated whether the edited models could recall the exact editing fact under editing prompts. The assessment was based on Efficacy Score (ES) representing the proportion of cases where $\mathbb{P}\left[o^{*}\right]>\mathbb{P}\left[o\right]$ [^4] Here, $\mathbb{P}\left[o\right]$ represents the probability that a model outputs $o$ given an editing prompt. after model editing, and Efficacy Magnitude (EM) representing the mean difference $\mathbb{P}\left[o^{*}\right]-\mathbb{P}\left[o\right]$.
  • $\bullet$
    Generalization verified whether the edited models could recall the editing fact under paraphrase prompts. It was measured by Paraphrase Score (PS) and Paraphrase Magnitude (PM), which were similar to ES and EM respectively.
  • $\bullet$
    Locality was to verify whether the output of the edited models for inputs out of editing scope remains unchanged after editing under neighborhood prompts. Different from efficacy, Neighborhood Score (NS) was to calculate the proportion of cases where $\mathbb{P}\left[o\right]>\mathbb{P}\left[o^{*}\right]$ after editing, and the Neighborhood Magnitude (NM) was to calculate the mean difference $\mathbb{P}\left[o\right]-\mathbb{P}\left[o^{*}\right]$.
  • $\bullet$
    Portability was to gauge the effectiveness of the edited models in transferring editing knowledge to related content. Given an additional triple $(o^{*},r^{\prime},o^{\prime})$, the relations $r$ and $r^{\prime}$ were combined to create a new relation $r^{\prime\prime}$. Then, the subject $s$ and relation $r^{\prime\prime}$ were formulated into a question where $o^{\prime}$ was the correct answer. This metric was to calculate the proportion of cases where the edited models could recall $o^{\prime}$.
  • $\bullet$
    Multi-hop shared the similar evaluation motivation as Portability, and was further extended to multi-hop questions entailed by the editing facts.
  • $\bullet$
    Reversibility was to evaluate the effectiveness of edited models in recalling the editing knowledge under reverse prompts. For QA form prompts, Reverse-QA Score (RQS) was utilized to calculate the proportion of cases where $\mathbb{P}\left[s\right]>\mathbb{P}\left[s^{*}\right]$ after model editing, and Reverse-QA Magnitude (RQM) was their mean difference. As for judgement form prompts, Reverse-Judgment Score (RJS) was defined to calculate the proportion of cases where $\mathbb{P}\left[yes\right]>\mathbb{P}\left[no\right]$ post editing, and Reverse-Judgment Magnitude (RJM) was their mean difference.
RS was denoted as the average of RQS and RJS [^5] Since there is no RQS in BAKE-J, only RJS is used to compute the harmonic mean.. Finally, the harmonic mean of ES, PS, NS, and RS was reported as Score (S) to comprehensively evaluate the performance of edited models in a more balanced manner.
5.3 Results and Comparison
Table 3 and  4 reported the knowledge editing results of different editing methods on BAKE-Q&J and BAKE-J respectively. The RQS and RJS results for each base LLM were 0, since the counterfactual editing triples that pre-existed in each LLM have been filtered out as mentioned in Section 4.2. Since Yao et al. (2023) and Zhong et al. (2023) evaluate editing only based on GPT-J, Table 5 reported the performance of editing methods on GPT-J in terms of portability, multi-hop, and reversibility.
Table 4: Evaluation results (%) of the BAKE-J dataset.
Editor Score Efficacy Generalization Locality Reversibility
$\mathrm{S}\uparrow$ $\mathrm{ES}\uparrow$ $\mathrm{EM}\uparrow$ $\mathrm{PS}\uparrow$ $\mathrm{PM}\uparrow$ $\mathrm{NS}\uparrow$ $\mathrm{NM}\uparrow$ $\mathrm{RJS}\uparrow$ $\mathrm{RJM}\uparrow$
GPT-2 XL (1.5B) 0.00 29.77 -6.28 31.04 -6.03 68.52 5.90 0.00 -0.02
FT 23.66 86.34 34.84 53.13 3.39 63.87 5.05 8.13 -0.02
KN 24.82 29.33 -4.47 29.52 -4.26 69.06 4.07 12.70 -0.02
MEND 0.28 91.27 33.59 73.89 15.33 28.45 -10.46 0.07 -0.02
MEMIT 19.21 82.88 42.58 68.48 16.77 66.31 5.05 6.01 -0.02
ROME 48.55 97.58 90.71 85.78 43.35 63.32 2.72 22.38 -0.01
BIRD $\textbf{51.99}^{\dagger}$ 97.14 90.86 86.51 44.07 63.09 2.61 $\textbf{25.49}^{\dagger}$0.01
GPT-J (6B) 0.00 22.80 -10.46 23.21 -10.57 74.00 10.12 0.00 -0.64
FT 23.34 83.23 35.58 45.79 1.12 71.09 9.44 8.10 -0.54
KN 39.56 38.10 -0.64 37.96 -0.51 58.24 0.44 31.90 0.02
MEND 0.75 92.09 51.99 69.23 19.38 34.29 -10.22 0.19 -0.65
MEMIT 34.83 95.67 91.08 84.43 46.85 69.43 6.97 12.84 -0.52
ROME 56.05 96.23 94.57 91.75 66.18 65.55 3.59 28.73 -0.32
BIRD $\textbf{57.95}^{\dagger}$ 96.26 94.61 91.9 67.55 65.33 3.42 $\textbf{30.82}^{\dagger}$0.30
LLaMA-1 (7B) 0.00 18.06 -16.69 18.32 -16.37 77.40 15.37 0.00 -1.68
FT 27.08 96.14 34.41 90.65 25.33 21.06 -13.58 12.70 -1.22
KN 26.94 18.85 -12.22 17.89 -11.67 76.49 10.43 37.80 -0.15
MEND 6.43 91.49 56.03 68.92 17.99 36.15 -8.19 1.76 -1.56
MEMIT 63.88 95.92 89.17 92.59 54.12 67.28 7.17 37.70 -0.45
ROME 64.71 95.70 91.60 90.05 46.25 69.28 9.12 38.62 -0.38
BIRD $\textbf{66.10}^{\dagger}$ 95.92 91.78 90.52 47.80 69.06 8.97 $\textbf{40.72}^{\dagger}$0.28
LLaMA-2 (7B) 0.00 15.99 -22.26 16.09 -22.30 80.11 20.92 0.00 -0.16
FT 30.00 95.32 30.14 89.92 24.19 23.39 -12.62 14.50 -0.25
KN 43.16 43.16 -17.58 42.81 -15.69 51.10 12.72 37.65 -0.05
MEND 12.53 88.51 54.10 64.30 17.00 44.06 -3.39 3.71 -0.31
MEMIT 63.64 96.11 90.84 90.76 57.81 70.41 12.06 36.73 -0.14
ROME 68.34 96.26 94.01 91.01 53.39 73.29 14.48 42.55 -0.06
BIRD $\textbf{69.47}^{\dagger}$ 96.75 93.92 91.16 50.79 73.03 14.75 $\textbf{44.30}^{\dagger}$0.04
Table 5: Comparison results (%) for portability, multi-hop and the proposed reversibility (in BAKE-Q&J) on GPT-J. Results except ours were cited from the published literature.
Editor Portability Multi-hop Reversibility
$\mathrm{RQS}$ $\mathrm{RJS}$
FT 6.29 1.9 0.68 6.90
KN 6.12 - 41.08 30.81
MEND 0.00 11.5 0.87 0.35
MEMIT 47.45 12.3 0.16 13.47
ROME 46.49 18.1 2.80 33.89
Evaluation in the Direction of Editing
All methods except KN performed well in terms of efficacy and generalization, showing that most of existing model editing methods are capable of effectively recalling the editing facts in the direction of editing. However, the performance in terms of locality is still limited and deserves further investigation. Moreover, with the increase in model size, the performance of a specific editing method continued to improve on these three metrics. For instance, the performance of ROME applied to LLaMA-2 was notably higher than that applied to GPT-2 XL in terms of ES, PS, and NS by 0.96%, 3.69%, and 11.96% respectively on the BAKE-Q&J dataset.
Evaluation in the Reverse Direction of Editing
Surprisingly, we observed that existing editing methods and LLMs failed catastrophically when evaluated in the reverse direction of editing, especially in the QA form. Taking LLaMA-2 edited by ROME on BAKE-Q&J as an example, it achieved 99.70% ES score which exhibited an impressive ability of recalling the editing facts in the direction of editing. However, it could only recall 0.26% (RQS) of the editing facts when evaluated in the reverse direction of editing using the QA format. Even when evaluated via the judgment form, it answered only 40.78% (RJS) of the questions correctly. These results show a surprising conclusion that, although current model editing methods and LLMs can effectively recall the editing facts in the direction of editing, they suffer serious deficiencies when evaluated in the reverse direction. Our results suggest that current editing methods are far from fully understanding the injected knowledge, but just simply memorize via hard encoding. These also echo the reversal curse in Berglund et al. (2023) and Grosse et al. (2023), and we hope to attract more attention to this difficult issue.
Comparison between Metrics
To provide an overview of various capabilities of existing editing methods, a comparison between reversibility and two follow-up metrics of portability (Yao et al., 2023) and multi-hop (Zhong et al., 2023) was conducted. As depicted in Table 5, the results of reversibility were significantly lower than those of portability across most of methods. Even compared to the multi-hop metric, the RQS results were generally lower. These results indicate that the proposed reversibility metric is indeed more challenging and points out the urgent shortcomings of current editing methods, requiring further research efforts.
There are two more findings: (1) The RJS results of GPT-2 XL and GPT-J edited by existing methods on BAKE-J (Table 4) were lower than those on BAKE-Q&J (Table 3), while other larger and more advanced LLMs were not like this. This might be that the relation categories in BAKE-J (many-to-one, many-to-many) were more difficult to make reverse judgments. Smaller models struggled with this difficulty, but as model sizes increased, their enhanced capabilities were enough to eliminate this problem. (2) In terms of reversibility, especially with regard to RJS, the performance of gradient-based methods (FT, MEND) is significantly lower than methods relying on locating knowledge neurons (KN, MEMIT, ROME). This indicates that gradient update exhibits a myopic nature, with a notably weak capability to update information in the reverse direction. Extensive experiments on more LLMs of different sizes will be part of our future work.
6 BIRD: Bidirectionally Inversible Relationship moDeling
The results in Section 5 demonstrate the ineffectiveness of existing model editing methods for knowledge update in the reverse direction of editing. In this section, we propose BIRD (Bidirectionally Inversible Relationship moDeling) to mitigate the reversal curse.
6.1 Our Method
Word2Vec (Mikolov et al., 2013) is proposed to use simple model architecture to train high quality word vectors, which can be performed with simple algebraic operations. For instance, $V(Paris)-V(France)\approx V(London)-V(England)$, where $V$ denotes the vector of the corresponding word. Here, we assume that this property in static word vectors still holds in the contextualized representations of language models. To extend to the form of factual triples, $R(Paris)-R(France)\approx R(the~{}capital~{}of)$ can be derived, where $R$ denotes the representations deduced from language models. Transposing terms to derive $R(Paris)\approx R(the~{}capital~{}of)+R(France)$, which can be further generalized to $R(object)\approx R(relation)+R(subject)$.
To make use of this property, this paper designs a set of simple yet effective objectives and incorporates them into Eq. (2) to improve ROME. The proposed objectives aim at incorporating the bidirectional relationships between subject and object in an edit fact into the updated model weights. To enhance the association of the new fact, the representation of the newly editing object is driven close to the representation of subject plus forward relation via the objective $\mathcal{L}_{1}(z)$ which is represented as:
$*\displaystyle\mathcal{L}_{1}(z)=H(s,r,o^{*}$ $,G(s),z)$ (4)
$=\frac{1}{N}\sum_{j=1}^{N}M$ $[\mathbb{R}_{G(m_{i^{\prime\prime}}^{(L)}:=z)}(s|x_{j}+s)$
$+\mathbb{R}(r|x_{j}+r)-\mathbb{R}(o^{*}|x_{j}+o^{*})].$
Here, $x_{j}$ is the sampled sentence mentioned in Section 3.2. $\mathbb{R}(s|x_{j}+s)$ means that given the representation of a prompt $x_{j}+s$ encoded by $G$, we choose the mean representation of subject $s$ from it, and similarly for others. $M$ is the $L2$ norm. $\mathcal{L}_{1}(z)$ uses the same operation in Eq. (2) to seek a vector z that minimize the $L2$ norm of the representation of subject plus forward relation minus the newly editing object. $G(s)$ denotes performing operation $G\left(m_{i^{\prime\prime}}^{\left(L\right)}:=z\right)$ exclusively on the last token $i^{\prime\prime}$ of subject $s$.
Reversely, the representation of subject is also driven close to the representation of the editing object plus backward relation via the objective $\mathcal{L}_{2}(z)$ which is represented as:
$\mathcal{L}_{2}(z)=H(o^{*},r^{-1}$ $,s,G(s),z)$ (5)
$=\frac{1}{N}\sum_{j=1}^{N}M$ $[\mathbb{R}(o^{*}|x_{j}+o^{*})+\mathbb{R}(r^{-1}|x_{j}+r^{-1})$
$-\mathbb{R}_{G(m_{i^{\prime\prime}}^{(L)}:=z)}(s|x_{j}+s)].$
On the other hand, to weaken the association of the old fact, the representation of the old object is driven away from the representation of subject along both the forward and backward relations, represented as $\mathcal{L}_{3}(z)$ = $H$($s$, $r$, $o$, $G(s)$, $z$) similarly to Eq. (4) and $\mathcal{L}_{4}(z)$=$H(o,r^{-1},s,G(s),z)$ similarly to Eq. (5) respectively.
Finally, these designed objectives are introduced and jointly optimized with the previous objective $\mathcal{L}(z)$ as:
$\mathcal{L}_{final}(z)$ $=\mathcal{L}(z)+\alpha[\mathcal{L}_{1}(z)+\mathcal{L}_{2}(z)-\beta(\mathcal{L}_{3}(z)+\mathcal{L}_{4}(z))],$ (6)
where $\alpha$ is the weight to balance the total contribution of the bidirectionally inversible relationship modeling and $\beta$ is utilized to control the degree to which the old facts are weakened.
6.2 Evaluation Results
As shown in Table 3 and  4, BIRD was also applied to four LLMs. Readers could refer to Appendix B.2 for the detailed hyperparameters for each model. In general, BIRD outperformed other editing methods in all experimental settings in terms of the overall Score. Particularly, BIRD improved the generalization ability when recalling the editing facts in the direction of editing given the same base model. More importantly, when evaluated in terms of reversibility, it can be seen that BIRD significantly outperformed other methods on RJS score across all settings, but the RQS performance was not as effective as RJS. From these experiments, we can draw the conclusion that BIRD can effectively improve the ability to distinguish, but it is still difficult for edited models to directly output the editing knowledge in the reverse direction of editing. Therefore, these results demonstrated the effectiveness of injecting the inversible relationships into the updated model weights bidirectionally, which mitigated the reversal curse to some extent. It also reveals that the task of bidirectional model editing is challenging and worth further investigation.
Figure 3: Average log probability of the original and desired answers after editing on the reverse-qa prompts. The closer the value is to 0, the higher the probability.
6.3 Probability of Desired Outputs
Although the evaluation results in Section 5.3 and  6.2 indicate that current editing methods are almost completely ineffective when evaluated in the QA form, we aim to further explore whether edited models increase the probability of desired outputs in the reverse direction of editing? To this end, the average log probability of the desired outputs before and after editing on LLaMA-2 (7B) on the reverse-qa prompts of BAKE-Q&J is illustrated in Figure 3. Readers can refer to Appendix C for the results of other LLMs.
Desired Answer After Editing
An effective editing method should increase the probability of the desired answer after editing. As shown in the orange columns in Figure 3, the probabilities kept unchanged or even became smaller compared with the base LLaMA-2, demonstrating these editing methods fail to increase the probability of the desired answer after editing.
Original Answer After Editing
An effective editing method should also reduce the probability of the original answer after editing. As shown in the blue columns in Figure 3, the probabilities became almost unchanged except for the KN method, which demonstrated the inadequacy of existing methods in reducing the probability of the original answer after editing.
Comparison between Original and Desired Answer After Editing
For all methods, the probability of the original answer was significantly higher than that of the desired answer after editing, and the margin between them was almost unchanged or became larger compared to the unedited LLaMA-2. This explained why the editing methods were almost ineffective in the QA form. Additionally, although the probability of the desired answer after editing decreased for the KN method, the margin with that of the original answer was minimized, which explained why the KN method achieved certain improvement in the QA form.
6.4 Effect of Objective Weights
Figure 4: Performance of BIRD with different $\alpha$ (a) or $\beta$ (b) on BAKE-J in terms of RJS.
To study the contribution of each component in Eq. (6), Figure 4 illustrated how the performance changed with respect to different hyperparameters $\alpha$ and $\beta$ on the BAKE-J dataset. When one coefficient was under evaluation, the other was assigned a default value. Due to computing limits, only 300 examples were selected.
It can be seen that the performance of BIRD was improved as $\alpha$ increased at the beginning, which shows the effectiveness of incorporating the bidirectionally inversible relationships between factual entities. Then, the performance dropped rapidly, as it should not be too large to affect the memorization of new facts. On the other hand, $\beta$ presented the roughly similar trend to $\alpha$, showing the necessity of weakening the association of the old fact.
7 Conclusion
We study bidirectional language model editing to assess if edited LLMs can recall the editing knowledge bidirectionally. A new evaluation metric of reversibility is introduced and a benchmark dubbed as BAKE is constructed. We find that current editing methods just simply memorize the newly injected knowledge via hard encoding, which are far from fully understanding. To mitigate the reversal curse, BIRD is proposed to enhance the association of new facts along both the forward and backward relations via a set of editing objectives. Experiments on four LLMs of different sizes demonstrate that the proposed editing method not only improves the generalization ability when recalling the editing facts in the direction of editing, but also achieves highly promising performance in the reverse direction of editing. We hope our work can facilitate future research on developing bidirectional language model editing.
Appendix A Bidirectional Handwritten Templates
Table 6 shows some examples of handwritten templates.
Table 6: The handwritten templates $T(r)$ for relation $r$ and the $T(r^{-1})$ for its inverse relation $r^{-1}$. Actually, there are several templates for each relation in both directions. Here we only display one template for each relation.
Relation (r) $T(r)$ $T(r^{-1})$
capital “The capital of {} is” “{} is the capital of”
chief executive officer “The chief executive officer of {} is” “{} is the chief executive officer of”
founded “{}, who has founded” “The {} has been founded by”
Appendix B Experimental Setup
B.1 Baseline Models
Five popular model editing methods were selected as baselines including:
  • $\bullet$
    FT (Zhu et al., 2020): this method simply performed gradient descent on the edits to update model parameters. It fine-tuned one layer in the model with a norm constraint on weight changes to prevent forgetfulness. Since the original authors did not publish their code, we followed Meng et al. (2022) re-implementation in their study.
  • $\bullet$
    KN (Dai et al., 2022) [^6] https://github.com/EleutherAI/knowledge-neurons.: it first selected neurons that were associated with knowledge expression via gradient-based attributions, and then modified MLP layer at the rows corresponding to those neurons by adding scaled embedding vectors.
  • $\bullet$
    MEND (Mitchell et al., 2022a) [^7] https://github.com/eric-mitchell/mend: it learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form.
  • $\bullet$
    ROME (Meng et al., 2022) [^8] https://github.com/kmeng01/rome: it first localized the factual knowledge at a specific layer in the transformer MLP modules, and then updated the knowledge by directly writing new key-value pairs in the MLP module.
  • $\bullet$
    MEMIT (Meng et al., 2023) [^9] https://github.com/kmeng01/memit: it extended ROME to edit a large set of facts and updated a sequence of MLP layers to update knowledge.
The ability of these methods were assessed based on EasyEdit (Wang et al., 2023), an easy-to-use knowledge editing framework which integrates the released codes and hyperparameters from previous methods.
B.2 Hyperparameters for BIRD
Table 7 shows the details of hyperparameters set for different LLMs.
Table 7: The hyperparameters for BIRD on each models.
Models (r) $\alpha$ $\beta$
GPT-2 XL 0.005 1.2
GPT-J 0.001 1.5
LLaMA-1 0.0005 0.8
LLaMA-2 0.002 0.8
Appendix C Log Probability for LLMs
Figure 3 depicts the log probability of the original and desired answers after editing on LLaMA-2 (7B). Figure 5,  6, and  7 show the log probability of other LLMs: GPT-2 XL (1.5B), GPT-J (6B), LLaMA-1 (7B), respectively.
Figure 5: Average log probability of the original and desired answers after editing on the reverse-qa prompts on GPT-2 XL (1.5B). The closer the value is to 0, the higher the probability.
Figure 6: Average log probability of the original and desired answers after editing on GPT-J (6B).
Figure 7: Average log probability of the original and desired answers after editing on LLaMA-1 (7B).