### v2

# Untying the Reversal Curse via Bidirectional Language Model Editing

Abstract
Large language models (LLMs) are prone to hallucinate unintended text due to false or outdated knowledge. Since retraining LLMs is resource intensive, there has been a growing interest in model editing. Despite the emergence of benchmarks and approaches, these unidirectional editing and evaluation have failed to explore the reversal curse. In this paper, we study bidirectional language model editing, aiming to provide rigorous evaluation to assess if edited LLMs can recall the editing knowledge bidirectionally. A metric of reversibility is introduced and a benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed to evaluate if post-edited models can recall the editing knowledge in the reverse direction of editing. We surprisingly observe that current editing methods and LLMs, while effective in recalling editing facts in the direction of editing, suffer serious deficiencies when evaluated in the reverse direction. To mitigate the reversal curse, a method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed, which designs a set of editing objectives to incorporate bidirectional relationships between subject and object into the updated model weights. Experiments show that BIRD improves the performance of four representative LLMs of different sizes via question answering and judgment.
Machine Learning, ICML
Figure 1: Comparison of unidirectional evaluation paradigms that assess whether edited models can recall the editing facts (a) via single-hop questions, or (b) via entailed questions in the direction of editing; and (c) the proposed bidirectional paradigm that assesses model editing in the reverse direction of editing. $f_{\theta}$ / $f_{\theta_{e}}$ denotes the models before / after editing.
1 Introduction
Large language models (LLMs) have shown impressive capabilities in understanding and generating text, improving performance on a variety of tasks (Chen et al., 2021; Ouyang et al., 2022; Zhang et al., 2023a). However, existing LLMs inevitably exhibit hallucinations due to false or outdated knowledge stored in their parameters (Zhang et al., 2023b; Peng et al., 2023; Ji et al., 2023). Considering the huge computing resource consumption of retraining LLMs and echoing green deep learning (Schwartz et al., 2020; Xu et al., 2021), there has been considerable attention on the research regarding knowledge editing (a.k.a., model editing(Sinitsin et al., 2020; Cao et al., 2021; Meng et al., 2022, 2023; Dai et al., 2022; Yao et al., 2023; Ma et al., 2024b), which aims at modifying knowledge by updating model weights.
Existing methods developed to modify the parameters of models can be generally categorized into two paradigms of meta-learning or locate-then-edit (Wang et al., 2023; Yao et al., 2023). The effectiveness of an editing method is usually evaluated along three dimensions of efficacy, generalization and locality. The ZsRE (Levy et al., 2017) and CounterFact (Meng et al., 2022) datasets are primarily utilized to assess whether edited models can recall the newly injected facts, as depicted in Figure 1(a). Besides, Yao et al. (2023) and Zhong et al. (2023) have introduced portability and MQuAKE respectively to extend the assessment of generalization by measuring whether edited models can correctly answer the questions entailed by the editing facts, as shown in Figure 1(b). Despite the emergence of benchmarks and techniques for editing, they all operate under the unidirectional paradigm following only the direction being edited. The reversal curse (Berglund et al., 2023) has not yet been explored in model editing, and it remains unclear whether edited models can reason and recall the editing facts in the reverse direction of editing, as shown in Figure 1(c).
In this paper, we study bidirectional language model editing, determining if edited LLMs can accurately recall the editing knowledge bidirectionally. Among various kinds of learned beliefs such as logical, spatial, or numerical knowledge, this paper investigates factual knowledge within LLMs, where each piece of knowledge is represented in the form of a (subject, relation, object) triple. A metric of reversibility is introduced to evaluate the effectiveness of edited models in recalling knowledge in the reverse direction of editing. A benchmark dubbed as Bidirectional Assessment for Knowledge Editing (BAKE) is constructed. When inverting the relation between subject and object, the taxonomy which categorizes relations between entities into four classes: one-to-one, one-to-many, many-to-one, many-to-many (Bordes et al., 2013), is considered. To characterize these symmetrical and asymmetrical relations, two evaluation forms of question answering (QA) and judgment are adopted for evaluation respectively.
Towards bidirectional model editing, a method named Bidirectionally Inversible Relationship moDeling (BIRD) is proposed to assist edited models to recall the editing facts in the reverse direction. A set of editing objectives is designed to incorporate the bidirectional and invertible relationships between subject and object into the updated model weights. Specifically, the representation of the newly editing object is driven close to the representation of the subject plus forward relation. Reversely, the representation of the subject is also driven close to the representation of the editing object plus the backward relation. Here, the relationships along both directions are inversible to each other. By this means, it can help characterize the bidirectional and inversible relationships between subject and object in factual knowledge, enabling models to establish factual associations bidirectionally during the editing process.
To establish the comparison of existing model editing methods on bidirectional model editing, this paper presents meticulous experiments involving as many methods as possible such as FT (Zhu et al., 2020), MEND (Mitchell et al., 2022), KN (Dai et al., 2022), ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023). Four representative LLMs of different sizes are selected as the base models including GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang & Komatsuzaki, 2021), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). We surprisingly observe that current editing methods and LLMs, while effective in recalling editing facts in the direction of editing, suffer serious deficiencies when evaluated in the reverse direction. Strikingly, the LLaMA-2 model edited by the state-of-the-art ROME (Meng et al., 2022) can recall 99.70% of the editing facts in the editing direction, but only 0.26% of the editing facts in the reverse direction. Besides, experiments demonstrate that the proposed editing method BIRD not only improves the generalization ability when recalling the editing facts in the direction of editing, but also achieves highly promising performance in the reverse direction.
In summary, our contributions are three-fold: (1) This study makes the first attempt to explore bidirectional language model editing. A metric of reversibility is introduced and a benchmark of BAKE is constructed to assess the reversibility of edited models. (2) This paper conducts meticulous experiments and presents surprising finding that existing methods and LLMs suffer from the severe reversal curse on model editing. (3) The BIRD method is proposed to mitigate the reversal curse in model editing, which incorporates the bidirectional and inversible relationships between subject and object into the updated model weights. All data and code will be available to facilitate reproducing our results. We hope that both our benchmark and method can shed light on bidirectional language model editing.
2 Related Work
Model Editing
Many studies have investigated model editing to modify the parameters of models. On the one hand, meta-learning methods train a hypernetwork to get gradient changes to update parameters (Cao et al., 2021; Mitchell et al., 2022). Cao et al. (2021) utilized a hypernetwork to predict parameter shift at test time. Mitchell et al. (2022) learned to transform the fine-tuning gradient into a low-rank decomposition of the gradient. On the other hand, locate-then-edit methods first locate knowledge neurons in LLMs that exhibit a positive correlation with a knowledge expression, and then modify them accordingly. In particular, Dai et al. (2022) computed the contribution of each neurons to a certain knowledge, then updated or erased knowledge by modifying these neurons with the embedding vectors of facts. Meng et al. (2022) located multi-layer perceptron (MLP) storing factual knowledge, and then edited such knowledge by injecting new key-value pair (Geva et al., 2021) in the MLP module. Besides, some benchmarks are proposed for assessing knowledge editing (Meng et al., 2022; Cohen et al., 2023; Ma et al., 2024a; Li et al., 2023; Gandikota et al., 2023). For example, Zhong et al. (2023) used the multi-hop questions to assess editing, suggesting that editing a particular fact implies that other facts need to be changed.
The Reversal Curse of LLMs
Recent works have shown a surprising failure of generalization in auto-regressive LLMs. Berglund et al. (2023) have verified that a model fine-tuned on a sentence of the form “A is B”, it will not automatically generalize to the reverse direction “B is A”. Moreover, the likelihood of the correct answer will not be higher than for a random answer. This is termed as the reversal curse. Contemporary to this work, Grosse et al. (2023) determined how much adding a given training example influences an LLM’s outputs. For instance, given the input A, what most influences the likelihood of B? In their experiments, training examples that match the order (“A precedes B”) are far more influential than examples with reverse order (“B precedes A”). Results show that if a pretrained model was not trained on facts in both directions, it would not generalize to both directions.
Compared with previous studies (Dai et al., 2022; Meng et al., 2022, 2023; Mitchell et al., 2022; Yao et al., 2023; Zhong et al., 2023) that are the most relevant to our work, a main difference should be highlighted. These approaches target only on unidirectional editing and evaluation of model editing, while this study explores bidirectional editing and evaluation. To the best of our knowledge, this paper makes the first attempt to introduce a metric of reversibility, build a benchmark for evaluating the reversibility of editing methods, and propose a simple yet effective method to mitigate the reversal curse in model editing.
3 Preliminary
3.1 Model Editing
The goal of model editing is to insert new facts into model parameters without retraining. In this paper, we study facts of the form ($s,r,o)$, consisting of a subject $s$, a relation $r$, and an object $o$ (e.g., $s$ = Eiffel Tower, $r$ = locate in, $o$ = Paris). Besides, we follow previous works (Zhong et al., 2023; Yao et al., 2023) to employ discrete prompts (question or cloze-style statement) to test whether a fact is in a language model. Editing a fact is to insert a new triple $(s,r,o^{*})$ in place of the current triple $(s,r,o)$, where these two triples share the same subject and relation. An editing operation is represented as $e=(s,r,o,o^{*})$ for brevity.
Given an edit $e$ and a model $f$ , model editing involves learning a function $K$ that yields an edited language model $f^{*}:K(f,e)=f^{*}$. To evaluate the effectiveness of various model editing methods, previous works focus on evaluation along three dimensions: efficacy, generalization and locality (Mitchell et al., 2022; Meng et al., 2022).
Efficacy validates whether the edited models could recall the exact editing fact when presented with the sole editing prompt $p$. The assessment is based on Efficacy Score (ES), which represents the new object $o^{*}$ exceeds $o$ in probability after model editing:  $\,\mathbbm{1}[P_{f^{*}}(o^{*}\,|\,p)>P_{f^{*}}(o\,|\,p)]\,$, where $\mathbbm{1}$ is the indicator function.
[…]
Locality [^1] Also referred to as specificity in Meng et al. (2022, 2023). is to verify whether the output of the edited models for inputs out of editing scope remains unchanged after editing under locality prompts. via Locality Score (LS): $\mathbb{E}_{p_{l}\in\mathcal{P}^{L}}[\mathbbm{1}[\,P_{f^{*}}(o_{l}\,|\,p_{l})>P_{f^{*}}(o^{*}\,|\,p_{l})]\,]$, where $o_{l}$ was the original answer of $p_{l}$.
For example, given an editing fact ($s$ = Eiffel Tower, $r$ = located in, $o$ = Paris, $o^{*}$=London), the editing prompt $p$ could be “Eiffel Tower is located in”, A paraphrase prompt could be “What is the location of the Eiffel Tower?”. A locality prompt and its original answer could be “Where is the Louvre?” and “Paris” respectively.
3.2 Rank-One Model Editing (ROME)
ROME (Meng et al., 2022) applies causal mediation analysis (Pearl, 2001; Vig et al., 2020) to locate the MLP modules that store facts and modify the second layers $W_{proj}^{l}$ in MLP. A new fact is represented as a key-value pair ($k_{*}$, $v_{*}$(Geva et al., 2021), which can be inserted into the MLP module of the model by solving a constrained least-squares problem (Bau et al., 2020). Denote $L$ as an edited MLP layer, $G$ as the model, and $e=(s,r,o,o^{*})$ as an edit. The key $k_{*}$ is obtained by calculating the representation of the last token of the subject $s$ outputted by $W_{fc}^{L}$.
The value $v_{*}$ encodes the new knowledge ($r$, $o^{*}$) as a property of $s$. To calculate it, ROME sets $v_{*}=\operatorname{argmin}_{z}\mathcal{L}(z)$, where the objective is denoted as:
$\mathcal{L}(z)=\frac{1}{N}\sum_{j=1}^{N}-\log\mathbb{P}_{G\left(m_{i}^{\left(L\right)}:=z\right)}\left[o^{*}\mid x_{j}+p\right]+$ (1)
$D_{\mathrm{KL}}\left(\mathbb{P}_{G\left(m_{i^{\prime}}^{\left(L\right)}:=z\right)}\left[x\mid p^{\prime}\right]\|\mathbb{P}_{G}\left[x\mid p^{\prime}\right]\right)\text{. }$
$x_{j}$ is the random sampled sequence. The first term in Eq. (1) seeks a vector z that, when substituted as the output of the MLP at token $i$ (notated $G\left(m_{i}^{\left(L\right)}:=z\right)$, and $i$ is the last subject token in $x_{j}+p$), will cause the model to predict the target object $o^{*}$ under the prompt $p$ (e.g., $p$=Eiffel Tower is located in, $o^{*}$=London). The second term minimizes the KL divergence (Ma et al., 2022) of predictions for the prompt $p^{\prime}$ (“{ subject } is”) to the unchanged model, preserving the model’s understanding of the subject’s essence ($i^{\prime}$ is the location of the last token of the subject in $p^{\prime}$).
Finally, the MLP weight $W_{proj}^{L}$ is updated with a rank-one update to insert the new fact:
$\hat{W}=W+\Lambda\left(C^{-1}k_{*}\right)^{T}\text{,}$ (2)
where $C$ is a constant by estimating the covariance of many existing keys. $\Lambda=\left(v_{*}-Wk_{*}\right)/\left(C^{-1}k_{*}\right)^{T}k_{*}$, representing the residual error of the new key–value pair on the original memory matrix.
Readers can refer to Appendix A for more details of ROME.
4 BAKE: Bidirectional Assessment for Knowledge Editing
Figure 2: Examples of four relation categories. A given relation is considered (a) one-to-one if a subject can be associated with at most one object; (b) one-to-many if a subject can be associated with multiple objects; (c) many-to-one if multiple subjects can be linked to the same object; or (d) many-to-many if multiple subjects can be linked to multiple objects.
The BAKE benchmark comprises two datasets of BAKE-Q&J and BAKE-J. Since it is difficult to directly measure the existing outdated and false knowledge in LLMs. Therefore, following previous works (Meng et al., 2022; Zhong et al., 2023; Yao et al., 2023), both datasets are designed for evaluating counterfactual edits in LLMs. When inverting the relation between subject and object, two evaluation forms of question answering (Q) and judgment (J) are adopted to characterize these symmetrical and asymmetrical relations.
4.1 Relation Category
There are various relations between entities in factual knowledge, which could be categorized into four classes: one-to-one, one-to-many, many-to-one, and many-to-many (Bordes et al., 2013), as shown in Figure 2.
For one-to-one and one-to-many relations, both question answering and judgment forms are employed since there is only one definite output after inverting the relation. Take ”developed” relation from one-to-many category in Figure 2(b) as an example, one can construct the factual statement “Apple developed iPhone”. When it is validated reversely, it could be presented to models in QA form like “iPhone is developed by ” or judgment form like “Whether iPhone is developed by Apple?” Similarly, one-to-one category shares the same property as one-to-many.
On the other hand, only the judgment form is employed for many-to-one and many-to-many relations, since there are alternative outputs to a given reversal relation. Suppose a “written by” relation from many-to-one category shown in Figure 2(c) is chosen to form a statement “The Trial is written by Kafka”. If presented to models for reverse validation in QA form like “Kafka has written ”, there may be many correct answers such as The Metamorphosis and The Castle. It is notable that edited models are expected to output The Trial intended for reversal evaluation instead of The Metamorphosis and The Castle. Therefore, evaluating the reversal relation for many-to-one and many-to-many categories via judgment is better. Similarly, the many-to-many category shares the same property as the many-to-one category.
4.2 Data Construction of BAKE-Q&J
This dataset is constructed based on Wikidata (Vrandecic & Krötzsch, 2014), a knowledge base containing millions of fact triples. First, we manually chose a total of 25 one-to-one and one-to-many relations, and all fact triples under each relation, denoted as $F(r)$= {$(s,r,o)$mid $s$ and $o$ are linked by $r$ in Wikidata}. For each relation $r$, ChatGPT (gpt-3.5-turbo) was used to generate some templates $T(r)$ and those $T(r^{-1})$ corresponding to its inverse relation [^2] Inverse relation $r^{-1}$ and its templates are only for evaluation. $r^{-1}$. Appendix B provides some examples of these templates. Entities can be replaced to form full prompts [^3] They could be in the form of question or cloze-style statement.: $\mathcal{P}(s,r)$ = {$t$.format($s$) $\mid$ $t\in T(r)$}, where .format() is string substitution. For example, a template for (r = written by) might be “{} is written by”, where “Hamlet” substitutes “{}”.
Constructing counterfactual edits
To construct an editing example, two triples sharing the relation $r$ are sampled from Wikidata, denoted as $(s,r,o)$ and $(s^{*},r,o^{*})$ respectively. An edit is represented as =$\mathcal{E}=\left\{s,r,o,o^{*},p\right\}$ to test efficacy, where the editing prompt $p\sim\mathcal{P}(s,r)$. Besides, to test generalization, a set of two semantically-equivalent paraphrase prompts $\mathcal{P}^{G}$ are sampled from $\mathcal{P}(s,r)\backslash\left\{p\right\}$. Moreover, to test locality, a set of triples that share the same object $o$ are collected: $\mathcal{S}=\left\{\left(s^{\prime},r^{\prime},o\right)\right\}$. For example, select ($s$= Eiffel Tower, $r$= located in, $o$= Paris), $\mathcal{S}$ might contain triple like ($s$= France, $r$= has capital, $o$= Paris). Then a set of prompts $\left\{\mathcal{P}\left(s^{\prime},r^{\prime}\right)\mid(s^{\prime},r^{\prime},o)\in\mathcal{S}\right\}$ are constructed to sample the locality prompts $\mathcal{P}^{L}$.
Constructing reverse prompts
As for the reversibility proposed in this paper, after modifying the knowledge from $(s,r,o)$ to $(s,r,o^{*})$, the edited model should be capable of deducing $(o^{*},r^{-1},s)$ if it is reversible. Following Yao et al. (2023) and  Zhong et al. (2023), we should ensure the triple $(o^{*},r^{-1},s)$ is not included (editing knowledge should not pre-exist) and $(o^{*},r^{-1},s^{*})$ is included (original knowledge should pre-exist) in the base LLM. If not, this editing example will be filtered out accordingly. Specifically, given $o^{*}$ and $r^{-1}$, we tested if a model can output $s^{*}$ rather than $s$, to deduce if the triple $(o^{*},r^{-1},s^{*})$ is included in the model. Furthermore, there are QA and judgment forms in this dataset and the reverse-qa prompt is represented as $\mathcal{P}^{Rq}=\left\{o^{*},r^{-1},s^{*},s,p^{-1}\right\}$, where $p^{-1}\sim\mathcal{P}(o^{*},r^{-1})$ and $s$ is the desired answer after editing (e.g., if $s$=Apple, $s^{*}$=Microsoft, $r$=developed, $o^{*}$=Windows, then $r^{-1}$=developed by, and $p^{-1}$ might be “Windows is developed by which company?”). Besides, the reverse-judge prompt $\mathcal{P}^{Rj}$ is a judgment question based on $p^{-1}$, for instance, “Whether Windows is developed by Apple?”, which should be answered with “yes” after editing or “no” before editing.
Table 1: An example in the BAKE-Q&J dataset. $a$ and $a^{*}$ in different columns represent the original answer and desired answer after editing respectively.
$\mathcal{E}$ $(s,r,o)$: (Paris, capital of, France)
$(s^{*},r,o^{*})$: (London, capital of, England)
(Paris, capital of, France $\rightarrow$ England)
$p$: Paris is the capital of ($a$: France, $a^{*}$: England)
$\mathcal{P}^{G}$ Paris, which is the capital of
Paris is the political center of which country?
($a$: France, $a^{*}$: England)
$\mathcal{P}^{L}$ Palace of Versailles and Louvre Museum are located in
Marseille is the city of which country?
($a$: France, $a^{*}$: France)
$\mathcal{P}^{Rq}$ What is the capital of England? ($a$: London, $a^{*}$: Paris)
$\mathcal{P}^{Rj}$ Whether the capital of England is Paris? ($a$: no, $a^{*}$: yes)
Table 2: The statistics of both BAKE-Q&J and BAKE-J.
Type $Num_{BAKE-Q\&J}$ $Num_{BAKE-J}$
Editing examples 4,694 3,184
Relations 25 20
Paraphrase prompts 8,811 5,917
Locality prompts 16,378 23,770
Reverse-qa prompts 4,694 -
Reverse-judge prompts 4,694 3,184
4.3 Data Construction of BAKE-J
The construction of this dataset is similar to BAKE-Q&J, and the difference lies in that 20 many-to-one and many-to-many relations in total were manually selected. Moreover, only the judgment form is adopted for validating the reversibility, which has been explained in Section 4.1.
4.4 Dataset Summary
Dataset Format
As shown in Table 1, each example in the BAKE benchmark is represented as a tuple ($\mathcal{E}$, $\mathcal{P}^{G}$, $\mathcal{P}^{L}$,$\mathcal{P}^{Rq}$,$\mathcal{P}^{Rj}$), where $\mathcal{E}$ is the editing knowledge which will be injected into the language model. $\mathcal{P}^{G}$ and $\mathcal{P}^{L}$ are the prompts utilized to validate generalization and locality respectively. $\mathcal{P}^{Rq}$ and $\mathcal{P}^{Rj}$ correspondingly represent the reverse-qa and reverse-judge prompts adopted to evaluate reversibility. $a$ and $a^{*}$ denote the original answer and desired answer after editing respectively.
Dataset Statistics
Table 2 summarizes the statistics of the BAKE-Q&J and BAKE-J datasets. To verify generalization and locality, there is at least one prompt for each editing example. Meanwhile, there is one reverse-qa prompt (only for BAKE-Q&J) and one reverse-judge prompt to verify reversibility. Both datasets consist of counterfactual edits utilized to investigate if edited models can recall the editing facts in both the editing and reverse directions.
5 BIRD: Bidirectionally Inversible Relationship moDeling
Figure 3: Illustration of BIRD which (a) enhances the association of the new fact bidirectionally, and (b) weakens the association of the old fact bidirectionally.
Word2Vec (Mikolov et al., 2013) is proposed to train word vectors with certain logical relationships, such as $V(Paris)-V(France)\approx V(London)-V(England)$, where $V$ denotes the vector of the corresponding word. Here, we assume that this property in static word vectors still holds in the contextualized representations of language models. To extend to the form of factual triples, $R(Paris)-R(France)\approx R(the~{}capital~{}of)$ can be derived, where $R$ denotes the representations deduced from language models. Transposing terms to derive $R(Paris)\approx R(the~{}capital~{}of)+R(France)$, which can be further generalized to $R(object)\approx R(relation)+R(subject)$.
To make use of this property, this paper designs a set of simple yet effective objectives and incorporates them into Eq. (1) to improve ROME. The proposed objectives aim at incorporating the bidirectional relationships between subject and object in an edit fact into the updated model weights. As shown in Figure 3 (a), to enhance the association of the new fact, the representation of the newly editing object is driven close to the representation of subject plus forward relation via the objective $\mathcal{L}_{1}(z)$ which is represented as:
$\mathcal{L}_{1}(z)=H(s,r,o^{*}$ $,G(s),z)$ (3)
$=\frac{1}{N}\sum_{j=1}^{N}M$ $[\mathbb{R}_{G(m_{i^{\prime\prime}}^{(L)}:=z)}(s|x_{j}+s)$
$+\mathbb{R}(r|x_{j}+r)-\mathbb{R}(o^{*}|x_{j}+o^{*})].$
Here, $x_{j}$ is the sampled sentence mentioned in Section 3.2. $\mathbb{R}(s|x_{j}+s)$ means that given the representation of a prompt $x_{j}+s$ encoded by $G$, we choose the mean representation of subject $s$ from it, and similarly for others. $M$ is the $L2$ norm. $\mathcal{L}_{1}(z)$ uses the same operation in Eq. (1) to seek a vector z that minimize the $L2$ norm of the representation of subject plus forward relation minus the newly editing object. $G(s)$ denotes performing operation $G\left(m_{i^{\prime\prime}}^{\left(L\right)}:=z\right)$ exclusively on the last token $i^{\prime\prime}$ of subject $s$.
Reversely, the representation of subject is also driven close to the representation of the editing object plus backward relation via the objective $\mathcal{L}_{2}(z)$ which is represented as:
$\mathcal{L}_{2}(z)=H(o^{*},r^{-1}$ $,s,G(s),z)$ (4)
$=\frac{1}{N}\sum_{j=1}^{N}M$ $[\mathbb{R}(o^{*}|x_{j}+o^{*})+\mathbb{R}(r^{-1}|x_{j}+r^{-1})$
$-\mathbb{R}_{G(m_{i^{\prime\prime}}^{(L)}:=z)}(s|x_{j}+s)].$
On the other hand, as shown in Figure 3 (b), to weaken the association of the old fact, the representation of the old object is driven away from the representation of subject along both the forward and backward relations, represented as $\mathcal{L}_{3}(z)$ = $H$($s$, $r$, $o$, $G(s)$, $z$) similarly to Eq. (3) and $\mathcal{L}_{4}(z)$=$H(o,r^{-1},s,G(s),z)$ similarly to Eq. (4).
Finally, these designed objectives are introduced and jointly optimized with the previous objective $\mathcal{L}(z)$ as:
$\mathcal{L}_{final}(z)$ $=\mathcal{L}(z)+\alpha[\mathcal{L}_{1}(z)+\mathcal{L}_{2}(z)-\beta(\mathcal{L}_{3}(z)+\mathcal{L}_{4}(z))],$ (5)
where $\alpha$ is to balance the total contribution of the bidirectionally inversible relationship modeling and $\beta$ is utilized to control the degree to which the old facts are weakened.
6 Experiments
6.1 Base LLMs and Editing Methods
Given limited computing resources, experiments were conducted on four LLMs of different sizes including GPT-2 XL (1.5B) (Radford et al., 2019), GPT-J (6B) (Wang & Komatsuzaki, 2021), LLaMA-1 (7B) (Touvron et al., 2023a), and LLaMA-2 (7B) (Touvron et al., 2023b). Readers can refer to the corresponding papers of each LLM for details. Five popular model editing methods were selected as the baselines including FT (Zhu et al., 2020), KN (Dai et al., 2022), MEND (Mitchell et al., 2022), ROME (Meng et al., 2022), and MEMIT (Meng et al., 2023). Readers can refer to Appendix C.1 for more details.
Table 3: Evaluation results (%) of the BAKE-Q&J and BAKE-J datasets. The settings for these methods were the same with the published literature. The row named after the model name represents the results evaluated on the base model. Numbers marked with $\dagger$ denoted that the improvement over the best performing baseline was statistically significant (t-test with $p$-value <$0.05$).
Editor BAKE-Q&J BAKE-J
Score Previous Reversibility Score Previous Reversibility
$\mathrm{S}\uparrow$ $\mathrm{ES}\uparrow$ $\mathrm{GS}\uparrow$ $\mathrm{LS}\uparrow$ $\mathrm{RQS}\uparrow$ $\mathrm{RJS}\uparrow$ $\mathrm{S}\uparrow$ $\mathrm{ES}\uparrow$ $\mathrm{GS}\uparrow$ $\mathrm{LS}\uparrow$ $\mathrm{RJS}\uparrow$
GPT-2 XL (1.5B) 0.00 29.12 29.57 70.15 0.00 0.00 0.00 29.77 31.04 68.52 0.00
FT 20.09 77.38 50.75 63.23 2.82 10.46 23.66 86.34 53.13 63.87 8.13
KN 27.41 30.77 29.7 70.68 14.42 16.05 24.82 29.33 29.52 69.06 12.70
MEND 3.89 93.29 77.89 27.82 1.82 0.25 0.28 91.27 73.89 28.45 0.07
MEMIT 16.54 73.43 62.77 64.31 0.29 9.88 19.21 82.88 68.48 66.31 6.01
ROME 41.55 98.74 91.05 52.34 3.44 32.25 48.55 97.58 85.78 63.32 22.38
BIRD $\textbf{42.68}^{\dagger}$ 98.7 91.32 52.19 2.91 $\textbf{34.50}^{\dagger}$ $\textbf{51.99}^{\dagger}$ 97.14 86.51 63.09 $\textbf{25.49}^{\dagger}$
GPT-J (6B) 0.00 22.92 23.00 77.15 0.00 0.00 24.44 22.80 23.21 74.00 0.00
FT 12.70 73.11 40.59 72.82 0.68 6.90 23.34 83.23 45.79 71.09 8.10
KN 41.77 38.99 38.57 61.09 41.08 30.81 39.56 38.10 37.96 58.24 31.90
MEND 2.36 94.29 72.50 35.91 0.87 0.35 0.75 92.09 69.23 34.29 0.19
MEMIT 21.73 96.21 86.99 64.59 0.16 13.47 34.83 95.67 84.43 69.43 12.84
ROME 42.53 98.01 94.60 53.27 2.80 33.89 56.05 96.23 91.75 65.55 28.73
BIRD 43.25 99.33 96.00 52.14 2.90 $\textbf{34.96}^{\dagger}$ $\textbf{57.95}^{\dagger}$ 96.26 91.9 65.33 $\textbf{30.82}^{\dagger}$
LLaMA-1 (7B) 0.00 19.41 19.99 80.96 0.00 0.00 0.00 18.06 18.32 77.40 0.00
FT 13.50 99.28 91.86 30.45 0.20 8.05 27.08 96.14 90.65 21.06 12.70
KN 29.23 28.21 29.06 75.35 10.58 26.67 26.94 18.85 17.89 76.49 37.80
MEND 5.12 75.00 60.62 47.23 0.46 2.28 6.43 91.49 68.92 36.15 1.76
MEMIT 40.95 99.28 95.34 58.38 0.15 33.19 63.88 95.92 92.59 67.28 37.70
ROME 41.79 99.19 94.13 62.11 0.20 33.75 64.71 95.70 90.05 69.28 38.62
BIRD $\textbf{43.61}^{\dagger}$ 99.33 95.67 61.95 0.10 $\textbf{36.22}^{\dagger}$ $\textbf{66.10}^{\dagger}$ 95.92 90.52 69.06 $\textbf{40.72}^{\dagger}$
LLaMA-2 (7B) 0.00 16.94 17.09 83.35 0.00 0.00 0.00 15.99 16.09 80.11 0.00
FT 13.78 97.72 91.37 32.66 0.21 8.18 30.00 95.32 89.92 23.39 14.50
KN 38.50 46.57 45.14 50.61 12.70 36.67 43.16 43.16 42.81 51.10 37.65
MEND 16.39 91.10 70.27 40.89 0.48 9.81 12.53 88.51 64.30 44.06 3.71
MEMIT 38.92 99.85 96.92 61.33 0.14 30.10 63.64 96.11 90.76 70.41 36.73
ROME 47.39 99.70 94.74 66.30 0.26 40.78 68.34 96.26 91.01 73.29 42.55
BIRD $\textbf{48.40}^{\dagger}$ 99.90 96.20 64.25 0.28 $\textbf{42.72}^{\dagger}$ $\textbf{69.47}^{\dagger}$ 96.75 91.16 73.03 $\textbf{44.30}^{\dagger}$
6.2 Evaluation Metrics
In addition to the proposed Reversibility, three basic metrics of Efficacy, Generalization and Locality (Meng et al., 2022, 2023) introduced in Section 3.1, as well as two follow-up metrics of Portability (Yao et al., 2023) and Multi-hop (Zhong et al., 2023) were adopted. Given two facts of $(s,r,o)$ and $(s^{*},r,o^{*})$, one editing operation updated $(s,r,o)$ to $(s,r,o^{*})$ and the edited model $f^{*}$ was obtained. And the reverse-qa prompts $\mathcal{P}^{Rq}$ were formulated by $o^{*}$ and $r^{-1}$, with answer $s$. Each metric was calculated as follows.
Portability was to gauge the effectiveness of the edited models in transferring editing knowledge to related content. Given an additional triple $(o^{*},r^{\prime},o^{\prime})$, the relations $r$ and $r^{\prime}$ were combined to create a new relation $r^{\prime\prime}$. Then, the subject $s$ and relation $r^{\prime\prime}$ were formulated into a question where $o^{\prime}$ was the correct answer. This metric was to calculate the proportion of cases where the edited models could recall $o^{\prime}$.
Multi-hop shared the similar evaluation motivation as Portability, and was further extended to multi-hop questions entailed by the editing facts.
[…]
For all the above metrics, the average results across all edits in each dataset were reported. Define RS was as the average of RQS and RJS [^4] Since there is no RQS in BAKE-J, only RJS is used to compute the harmonic mean., the harmonic mean of ES, GS, LS, and RS was reported as Score (S) to comprehensively evaluate the performance of edited models in a more balanced manner.
6.3 Results and Comparison
Table 3 reported the results of different methods on BAKE-Q&J and BAKE-J. The RQS and RJS results for each base LLM were 0, since the counterfactual editing triples that pre-existed in each LLM have been filtered out as mentioned in Section 4.2. Since Yao et al. (2023) and Zhong et al. (2023) evaluated editing only based on GPT-J, Table 4 reported the editing performance on GPT-J in terms of portability, multi-hop, and reversibility.
Table 4: Comparison results (%) for portability, multi-hop and the proposed reversibility (in BAKE-Q&J) on GPT-J. Results except ours were cited from the published literature.
Editor Portability Multi-hop Reversibility
$\mathrm{RQS}$ $\mathrm{RJS}$
FT 6.29 1.9 0.68 6.90
KN 6.12 - 41.08 30.81
MEND 0.00 11.5 0.87 0.35
MEMIT 47.45 12.3 0.16 13.47
ROME 46.49 18.1 2.80 33.89
Evaluation in the Direction of Editing
All methods except KN performed well in terms of efficacy and generalization, showing that most of existing editing methods are capable of effectively recalling the editing facts in the direction of editing. However, the performance in terms of locality is still limited and deserves further investigation. Moreover, with the increase in model size, the performance of a specific editing method continued to improve on these three metrics. For instance, the performance of ROME applied to LLaMA-2 was notably higher than that applied to GPT-2 XL in terms of ES, GS, and LS by 0.96%, 3.69%, and 11.96% respectively on the BAKE-Q&J dataset.
Evaluation in the Reverse Direction of Editing
Surprisingly, we observed that existing editing methods and LLMs failed catastrophically when evaluated in the reverse direction of editing, especially in the QA form. Taking LLaMA-2 edited by ROME on BAKE-Q&J as an example, it achieved 99.70% ES score which exhibited an impressive ability of recalling the editing facts in the direction of editing. However, it could only recall 0.26% (RQS) of the editing facts when evaluated in the reverse direction of editing using the QA format. Even when evaluated via the judgment form, it answered only 40.78% (RJS) of the questions correctly. These results show a surprising conclusion that, although current model editing methods and LLMs can effectively recall the editing facts in the direction of editing, they suffer serious deficiencies when evaluated in the reverse direction. Our results suggest that current editing methods are far from fully understanding the injected knowledge, but just simply memorize via hard encoding. These also echo the reversal curse in Berglund et al. (2023) and Grosse et al. (2023), and we hope to attract more attention to this difficult issue.
Comparison between Metrics
To provide an overview of various capabilities of existing editing methods, a comparison between reversibility and two follow-up metrics of portability (Yao et al., 2023) and multi-hop (Zhong et al., 2023) was conducted. As depicted in Table 4, the results of reversibility were significantly lower than those of portability across most methods. Even compared to the multi-hop metric, the RQS results were generally lower. These indicate that the proposed reversibility is indeed more challenging and point out the urgent shortcomings of current editing methods, requiring further efforts.
Another notable observation pertains to reversibility, particularly in the context of RJS. The performance of gradient-based methods (FT, MEND) is markedly lower compared to methods that rely on locating knowledge neurons (KN, MEMIT, ROME). This suggests that gradient updates demonstrate a myopic nature, showing a notably weak capacity to update information in the reverse direction.
6.4 Results of BIRD
As shown in Table 3, BIRD was also applied to four LLMs. Readers could refer to Appendix C.2 for the detailed hyperparameters for each model. In general, BIRD outperformed other editing methods in all experimental settings in terms of the overall Score. Particularly, BIRD improved the generalization ability when recalling the editing facts in the direction of editing given the same base model. More importantly, when evaluated in terms of reversibility, it can be seen that BIRD significantly outperformed other methods on RJS score across all settings, but the RQS performance was not as effective as RJS. From these experiments, we can draw the conclusion that BIRD can effectively improve the ability to distinguish, but it is still difficult for edited models to directly output the editing knowledge in the reverse direction of editing. Therefore, these results demonstrated the effectiveness of injecting the inversible relationships into the updated model weights bidirectionally, which mitigated the reversal curse to some extent. It also reveals that the task of bidirectional model editing is challenging and worth further investigation.
Figure 4: Average log probability of the original and desired answers after editing on the reverse-qa prompts. The closer the value is to 0, the higher the probability.
6.5 Probability of Desired Outputs
Although the evaluation results in Section 6.3 and  6.4 indicate that current editing methods are almost completely ineffective when evaluated in the QA form, we aim to further explore whether edited models increase the probability of desired outputs in the reverse direction of editing? To this end, the average log probability of the desired outputs before and after editing on LLaMA-2 (7B) on the reverse-qa prompts of BAKE-Q&J is illustrated in Figure 4. Readers can refer to Appendix D for the results of other LLMs.
Desired Answer After Editing. An effective editing method should increase the probability of the desired answer. The orange columns in Figure 4 show that the probabilities kept unchanged or even became smaller compared with the base LLaMA-2, demonstrating these methods fail to increase the probability of the desired answer.
Original Answer After Editing An effective editing method should also reduce the probability of the original answer. As shown in the blue columns in Figure 4, the probabilities became almost unchanged except for the KN method, which demonstrated the inadequacy of existing methods in reducing the probability of the original answer.
Comparison between Original and Desired Answer After Editing For all methods, the probability of the original answer was significantly higher than that of the desired answer after editing, and the margin between them was almost unchanged or became larger compared to the unedited LLaMA-2. This explained why the editing methods were almost ineffective in the QA form. Additionally, although the probability of the desired answer after editing decreased for the KN method, the margin with that of the original answer was minimized, which explained why the KN method achieved certain improvement in the QA form.
6.6 Effect of Objective Weights
Figure 5: Performance of BIRD with different $\alpha$ (a) or $\beta$ (b) on BAKE-J in terms of RJS.
To study the contribution of each component in Eq. (5), Figure 5 illustrated how the performance changed with respect to different hyperparameters $\alpha$ and $\beta$ on the BAKE-J dataset. When one coefficient was under evaluation, the other was assigned a default value.
It can be seen that the performance of BIRD was improved as $\alpha$ increased at the beginning, highlighting the effectiveness of incorporating the bidirectionally inversible relationships between factual entities. Subsequently, the performance dropped rapidly, as it should not be too large to affect the memorization of new facts. On the other hand, $\beta$ presented the roughly similar trend to $\alpha$, underscoring the necessity of weakening the association of the old fact.
7 Conclusion
We study bidirectional language model editing to assess if edited LLMs can recall the editing knowledge bidirectionally. A metric of reversibility is introduced and a benchmark dubbed as BAKE is constructed. We find that current editing methods just simply memorize the newly injected knowledge via hard encoding, which are far from fully understanding. To mitigate the reversal curse, BIRD is proposed to enhance the association of new facts along both the forward and backward relations via a set of editing objectives. Experiments on four LLMs of different sizes demonstrate that the proposed editing method not only improves the generalization ability when recalling the editing facts in the direction of editing, but also achieves highly promising performance in the reverse direction of editing. We hope our work can facilitate future research on developing bidirectional language model editing.
There are several limitations to our work. Firstly, although the performance of RJS is improved by BIRD, it still does not perform well in RQS. Developing more efficient bidirectional editing methods is an important direction for future research. Secondly, this work focuses on factual knowledge, but we have not investigated other types of learned beliefs, such as logical, spatial, or numerical knowledge. Thirdly, our research focuses on autoregressive models, and further exploration is needed for non-autoregressive models. Finally, this work uses editing methods that modifying parameters of models, while the performance of methods that preserve model parameters remains to be explored.
Impact Statements
The research centers on a pivotal facet of LLMs: mitigating hallucinations through model editing, which aim to facilitate better human-AI interactions and contribute to the development of safer and more responsible AI technologies. This not only enhances the scientific community’s comprehension of the strengths and limitations inherent in existing editing methods but also brings to light the challenges associated with the reversal curse in language model editing. The acquired insights from our research contribute to the current understanding and serve as a catalyst for future investigations into bidirectional relationships within language models. Furthermore, our work highlights the importance of responsible and transparent approaches, encouraging researchers to actively contribute to the ongoing discourse on responsible AI. This emphasis is pivotal in enhancing the overall reliability and trustworthiness of language models in real-world applications.
Appendix A Rank-One Model Editing (ROME)
ROME (Meng et al., 2022) applies causal mediation analysis (Pearl, 2001; Vig et al., 2020) to locate the MLP modules that store facts. An MLP module consists of two layers, where the first and second layers are denoted as $W_{fc}^{l}$ and $W_{proj}^{l}$ respectively. $W_{proj}^{l}$ is considered as a linear associative memory and an editing area. A new fact is represented as a key-value pair ($k_{*}$, $v_{*}$(Geva et al., 2021), which can be inserted into the MLP module of the model by solving a constrained least-squares problem (Bau et al., 2020). All that remains is to choose the appropriate $k_{*}$ and $v_{*}$.
Denote $L$ as an edited MLP layer, $G$ as the model, and $e=(s,r,o,o^{*})$ as an editing operation. The key $k_{*}$ is obtained by calculating the average representation of the last token of the subject $s$ outputted by $W_{fc}^{L}$ as follows:
$k_{*}$ $=\frac{1}{N}\sum_{j=1}^{N}k\left(x_{j}+s\right)\text{,}$ (6)
where $\text{where }k(x)$ $=\sigma\left(W_{fc}^{\left(L\right)}\gamma\left(a_{[x],i}^{\left(L\right)}+h_{[x],i}^{\left(L-1\right)}\right)\right)\text{. }$
Here, $x_{j}$ is the random sampled sequence and $i$ is the location of the last token of the subject in sentence $x$. $h$ is the hidden states, $a$ is the attention, $\gamma$ is the layernorm, and $\sigma$ is the activation function.
The value $v_{*}$ encodes the new knowledge ($r$, $o^{*}$) as a property of $s$. To calculate it, ROME sets $v_{*}=\operatorname{argmin}_{z}\mathcal{L}(z)$, where the objective is denoted as:
$\mathcal{L}(z)=\frac{1}{N}\sum_{j=1}^{N}-\log\mathbb{P}_{G\left(m_{i}^{\left(L\right)}:=z\right)}\left[o^{*}\mid x_{j}+p\right]+$ (7)
$D_{\mathrm{KL}}\left(\mathbb{P}_{G\left(m_{i^{\prime}}^{\left(L\right)}:=z\right)}\left[x\mid p^{\prime}\right]\|\mathbb{P}_{G}\left[x\mid p^{\prime}\right]\right)\text{. }$
The first term in Eq. (7) seeks a vector z that, when substituted as the output of the MLP at the token $i$ (notated $G\left(m_{i}^{\left(L\right)}:=z\right)$ and here $i$ is the end of the subject in prompt $x_{j}+p$), will cause the network to predict the target object $o^{*}$ in response to the factual prompt $p$ (e.g., $p$=Eiffel Tower is located in, $o^{*}$=London). The second term minimizes the KL divergence (Ma et al., 2022) of predictions for the prompt $p^{\prime}$ (of the form ” { subject } is a”) to the unchanged model, which helps preserve the model’s understanding of the subject’s essence ($i^{\prime}$ is the location of the last token of the subject in $p^{\prime}$).
Finally, the MLP weight $W_{proj}^{L}$ is updated with a rank-one update to insert the new fact:
$\hat{W}=W+\Lambda\left(C^{-1}k_{*}\right)^{T}\text{,}$ (8)
where $C$ is a constant by estimating the covariance of many existing keys. $\Lambda=\left(v_{*}-Wk_{*}\right)/\left(C^{-1}k_{*}\right)^{T}k_{*}$, representing the residual error of the new key–value pair on the original memory matrix.
Readers can refer to Meng et al. (2022) for more details of ROME.
Appendix B Bidirectional Templates
Table 5 shows some examples of templates.
Table 5: The templates $T(r)$ for relation $r$ and the $T(r^{-1})$ for its inverse relation $r^{-1}$. Actually, there are several templates for each relation in both directions. Here we only display one template for each relation.
Relation (r) $T(r)$ $T(r^{-1})$
capital “The capital of {} is” “{} is the capital of”
chief executive officer “The chief executive officer of {} is” “{} is the chief executive officer of”
founded “{}, who has founded” “The {} has been founded by”
currency “{}, which has the currency” “The {} is the currency utilized in”
Appendix C Experimental Setup
C.1 Baseline Models
Five popular model editing methods were selected as baselines including:
  • $\bullet$
    FT (Zhu et al., 2020): this method simply performed gradient descent on the edits to update model parameters. It fine-tuned one layer in the model with a norm constraint on weight changes to prevent forgetfulness. Since the original authors did not publish their code, we followed Meng et al. (2022) re-implementation in their study.
  • $\bullet$
    KN (Dai et al., 2022) [^5] https://github.com/EleutherAI/knowledge-neurons.: it first selected neurons that were associated with knowledge expression via gradient-based attributions, and then modified MLP layer at the rows corresponding to those neurons by adding scaled embedding vectors.
  • $\bullet$
    MEND (Mitchell et al., 2022) [^6] https://github.com/eric-mitchell/mend: it learned a hypernetwork to produce weight updates by decomposing the fine-tuning gradients into rank-1 form.
  • $\bullet$
    ROME (Meng et al., 2022) [^7] https://github.com/kmeng01/rome: it first localized the factual knowledge at a specific layer in the transformer MLP modules, and then updated the knowledge by directly writing new key-value pairs in the MLP module.
  • $\bullet$
    MEMIT (Meng et al., 2023) [^8] https://github.com/kmeng01/memit: it extended ROME to edit a large set of facts and updated a sequence of MLP layers to update knowledge.
The ability of these methods were assessed based on EasyEdit (Wang et al., 2023), an easy-to-use knowledge editing framework which integrates the released codes and hyperparameters from previous methods.
C.2 Hyperparameters for BIRD
Table 6 shows the details of hyperparameters set for different LLMs.
Table 6: The hyperparameters for BIRD on each models.
Models (r) $\alpha$ $\beta$
GPT-2 XL 0.005 1.2
GPT-J 0.001 1.5
LLaMA-1 0.0005 0.8
LLaMA-2 0.002 0.8
Appendix D Log Probability for LLMs
Figure 4 depicts the log probability of the original and desired answers after editing on LLaMA-2 (7B). Figure 6,  7, and  8 show the log probability of other LLMs: GPT-2 XL (1.5B), GPT-J (6B), LLaMA-1 (7B), respectively.
Figure 6: Average log probability of the original and desired answers after editing on the reverse-qa prompts on GPT-2 XL (1.5B). The closer the value is to 0, the higher the probability.
Figure 7: Average log probability of the original and desired answers after editing on GPT-J (6B).
Figure 8: Average log probability of the original and desired answers after editing on LLaMA-1 (7B).